# G2D2: Gradient-guided Discrete Diffusion for image inverse problem solving

### v1

@@ Appendix B Proofs
    $\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[-\log q(\mathbf{y}|\mathbf{z}_{0})\right]=-\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\log q(\mathbf{y}|\mathbf{z}_{0}).$ (27)
Thus, the KL divergence between $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and $q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ can be decomposed as follows:
$D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t})\right)-\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[\log q(\mathbf{y}|\mathbf{z}_{0})\right].$ (28)
@@
Lemma B.3.
The marginal distribution $q_{\text{sampling}}(\mathbf{z}_{0}|\mathbf{y})$ is identical to the target distribution $q(\mathbf{z}_{0}|\mathbf{y})$.
@@
We aim to show that:
$q_{\text{sampling}}(\mathbf{z}_{0}|\mathbf{y})=q(\mathbf{z}_{0}|\mathbf{y}).$ (29)
Starting from the definition of $q_{\text{sampling}}$, we have:
$q_{\text{sampling}}(\mathbf{z}_{0}|\mathbf{y})=\sum_{\mathbf{z}_{1:T}}q_{\text{sampling}}(\mathbf{z}_{0:T}|\mathbf{y})=\sum_{\mathbf{z}_{1:T}}q(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y}).$ (30)
Next, we can express $q(\mathbf{z}_{T}|\mathbf{y})$ using the sum over $\mathbf{z}^{\prime}_{0}$:
$q(\mathbf{z}_{T}|\mathbf{y})=\sum_{\mathbf{z}^{\prime}_{0}}q(\mathbf{z}^{\prime}_{0}|\mathbf{y})q(\mathbf{z}_{T}|\mathbf{z}^{\prime}_{0}).$ (31)
Substituting Eq. (31) into Eq. (30), we get:
$q_{\text{sampling}}(\mathbf{z}_{0}|\mathbf{y})=\sum_{\mathbf{z}_{1:T}}\sum_{\mathbf{z}^{\prime}_{0}}q(\mathbf{z}^{\prime}_{0}|\mathbf{y})q(\mathbf{z}_{T}|\mathbf{z}^{\prime}_{0})\prod_{t=1}^{T}q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y}).$ (32)
Rewriting the sums, we have:
$q_{\text{sampling}}(\mathbf{z}_{0}|\mathbf{y})=\sum_{\mathbf{z}^{\prime}_{0}}q(\mathbf{z}^{\prime}_{0}|\mathbf{y})\left(\sum_{\mathbf{z}_{1:T}}q(\mathbf{z}_{T}|\mathbf{z}^{\prime}_{0})\prod_{t=1}^{T}q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right).$ (33)
The inner sum over $\mathbf{z}_{1:T}$ represents the probability of transitioning from $\mathbf{z}^{\prime}_{0}$ to $\mathbf{z}_{0}$ through the forward and reverse processes. In the star-shaped noise process, the forward transition $q(\mathbf{z}_{T}|\mathbf{z}^{\prime}_{0})$ followed by the reverse transitions $\prod_{t=1}^{T}q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ effectively collapses back to $\mathbf{z}^{\prime}_{0}$, since the reverse transitions are designed to invert the forward process.
Therefore, the inner sum simplifies to:
$\sum_{\mathbf{z}_{1:T}}q(\mathbf{z}_{T}|\mathbf{z}^{\prime}_{0})\prod_{t=1}^{T}q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})=\delta_{\mathbf{z}_{0},\mathbf{z}^{\prime}_{0}},$ (34)
where $\delta$ is the Kronecker delta function, equal to 1 when $\mathbf{z}_{0}=\mathbf{z}^{\prime}_{0}$ and 0 otherwise.
Thus, we have:
$q_{\text{sampling}}(\mathbf{z}_{0}|\mathbf{y})=\sum_{\mathbf{z}^{\prime}_{0}}q(\mathbf{z}^{\prime}_{0}|\mathbf{y})\delta_{\mathbf{z}_{0},\mathbf{z}^{\prime}_{0}}=q(\mathbf{z}_{0}|\mathbf{y}),$ (35)
which proves the lemma. ∎
Appendix C Details on Experiments
C.1 Image inverse problems
@@
In our image inverse problem experiments, the definition and implementation of the forward operator are based on the DPS implementation [^2] https://github.com/DPS2022/diffusion-posterior-sampling. To ensure a diverse representation of ImageNet classes without genre bias, we select a subset consisting of 100 images from classes $0,10,\ldots,990$ using the imagenet_val_1k.txt provided by Pan et al. (2021) [^3] https://github.com/XingangPan/deep-generative-prior/. For our experiments with the FFHQ dataset, we use images $0,1,\ldots,99$ from the validation set.
C.2 Implementation Details of G2D2 in Inverse Problem Settings
@@
In the continuous optimization phase, we optimize the parameters $\bm{\alpha}$ of the categorical distribution using the Adam optimizer. The optimization objective is a weighted sum of the KL divergence term and the likelihood term, defined as:
@@
The marginalization over $\mathbf{z}_{0}$ in line 10 of Algorithm 1, specifically the term $\sum_{\mathbf{z}_{0}}q(\mathbf{z}_{t-1}|\mathbf{z}_{0})\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$, can be computed in closed form. This computation is feasible because both distributions involved in the marginalization are dimensionally independent categorical distributions, as discussed by Austin et al. (2021) and Gu et al. (2022).
@@
Some parameters are dynamically adjusted during inference. Both the learning rate for Adam ($l_{\text{Adam}}$) and the KL divergence coefficient ($\eta_{\text{KL}}$) are scheduled using weight vectors that decay logarithmically over the inference steps. These weights are computed based on initial scaling factors.
@@
$l_{\text{Adam}}(t)=l_{\text{Adam, base}}\cdot w_{\text{lr}}(t),\quad\eta_{\text{KL}}(t)=\eta_{\text{KL, base}}\cdot w_{\text{KL}}(t).$
@@
$l_{\text{Adam, base}}$ 1.0
$\lambda_{\text{lr, schedule}}$ 0.5
ImageNet Super-resolution $\eta_{\text{KL, base}}$ 0.0001
@@
$l_{\text{Adam, base}}$ 0.5
$\lambda_{\text{lr, schedule}}$ 2.0
@@
$l_{\text{Adam, base}}$ 0.5
$\lambda_{\text{lr, schedule}}$ 0.5
@@
$l_{\text{Adam, base}}$ 0.3
@@
C.3 G2D2 with Markov Noise Process
@@
5:         Initialize: $\bm{\alpha}_{t}=\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$
@@
7:         Initialize: $\bm{\alpha}_{t}=\exp(\gamma\log\bm{\alpha}_{t+1}+(1-\gamma)\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t}))$
@@
C.4 Settings for comparison methods
In this subsection, we detail the experimental settings for the comparison method.
@@ C.4 Settings for comparison methods
(Chung et al., 2023) We use the same parameter settings as described in the original paper. The guidance scale is set to 1.0 for FFHQ & Super resolution, 1.0 for FFHQ & Gaussian deblurring, 1.0 for ImageNet & Super resolution, and 0.4 for ImageNet & Gaussian deblurring. The number of time steps is set to 1000. For pre-trained models, we use the unconditional model provided by Dhariwal & Nichol (2021)  [^5] https://github.com/openai/guided-diffusion for ImageNet. For FFHQ, we use the model provided by Choi et al. (2021)  [^6] https://github.com/jychoi118/ilvr_adm.
@@
(Rout et al., 2023) We use the official implementation  [^8] https://github.com/LituRout/PSLD. For the pre-trained model, we employ stable-diffusion v-1.5 (Rombach et al., 2022)  [^9] https://github.com/CompVis/stable-diffusion. As this model handles 512$\times$512 pixel images, we first upscale the ground truth image to 512$\times$512. We then apply the forward operator to the upscaled image and use the result as observed data for our method. Finally, we downsample the output to 256$\times$256. For hyperparameters, we use $\eta=1.0$ and $\gamma=0.1$.
@@
C.5 GPU memory usage and computational speed
We analyze the GPU memory consumption and computational speed of our proposed method, G2D2, in comparison with other methods. Table 4 presents a overview of these metrics for various methods. The measurement are conducted using a single NVIDIA A6000 GPU for the Gaussian deblurring task on ImageNet. G2D2 has the lowest memory usage among all methods and the fastest computational speed among gradient-based methods.
@@ C.5 GPU memory usage and computational speed
Method GPU Memory Usage (GiB) Computational Time (s)
@@
C.6 Impact of the Forget Coefficient
@@
C.7 Additional qualitative results of G2D2 and comparison methods.

### v2

@@ Appendix D Proofs
    $\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[-\log q_{\text{star}}(\mathbf{y}|\mathbf{z}_{0})\right]=-\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\log q_{\text{star}}(\mathbf{y}|\mathbf{z}_{0}).$ (26)
Thus, the KL divergence between $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ can be decomposed as follows:
$D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})\right)-\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[\log q_{\text{star}}(\mathbf{y}|\mathbf{z}_{0})\right].$ (27)
@@
{restatable}
lemmaequimarginal The marginal distribution $q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})$ is identical to the target distribution $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{y})$.
@@
We aim to show that
$q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})=\;q_{\text{star}}(\mathbf{z}_{0}|\mathbf{y}).$ (28)
Recall first that $q_{\text{star}}(\mathbf{z}_{0:T}|\mathbf{y})$ is our original “star-shaped” joint posterior, and we define its “star-decomposed” variant by
$q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})=\;q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y}).$ (29)
By the definition of marginalization, to get $q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})$, we integrate over all the latent variables $\mathbf{z}_{1},\dots,\mathbf{z}_{T}$:
$q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})$ $=\;\sum_{\mathbf{z}_{1:T}}q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y}).$ (30)
Substituting our definition of $q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})$ into the sum yields
$q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})$ $=\;\sum_{\mathbf{z}_{1:T}}\Bigl[q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\Bigr].$ (31)
Here, $q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})$ and each factor $q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ are just conditional distributions of the star-shaped model.
Next, we marginalize over $\mathbf{z}_{T}$, then $\mathbf{z}_{T-1}$, and so on, one index at a time. That is, we view
$\sum_{\mathbf{z}_{T}}q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\,q_{\text{star}}(\mathbf{z}_{T-1}|\mathbf{z}_{T},\mathbf{y})=q_{\text{star}}(\mathbf{z}_{T-1}|\mathbf{y}),$
because summation over $\mathbf{z}_{T}$ of $q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\,q_{\text{star}}(\mathbf{z}_{T-1}|\mathbf{z}_{T},\mathbf{y})$ collapses exactly to the marginal $q_{\text{star}}(\mathbf{z}_{T-1}|\mathbf{y})$. This step follows directly from the chain rule of probability (or the law of total probability).
We then apply the same idea repeatedly: first summing out $\mathbf{z}_{T}$, which gives us a factor depending only on $\mathbf{z}_{T-1}$; then summing out $\mathbf{z}_{T-1}$ to obtain a factor depending on $\mathbf{z}_{T-2}$; and so on, down through $\mathbf{z}_{1}$. Performing these consecutive sums from $t=T$ down to $t=1$ eventually leaves only $\mathbf{z}_{0}$ in the expression. Hence, we get
$q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})=\;q_{\text{star}}(\mathbf{z}_{0}|\mathbf{y}),$ (32)
which shows that both star-decomposed and original star-shaped posteriors yield the same marginal distribution over $\mathbf{z}_{0}$. ∎
Appendix E Details on Experiments
E.1 Image inverse problems
@@
In our image inverse problem experiments, the definition and implementation of the forward operator are based on the DPS implementation [^2] https://github.com/DPS2022/diffusion-posterior-sampling. To ensure a diverse representation of ImageNet classes without genre bias, we select 1000 images covering all classes $0,1,2,\ldots,999$ using the imagenet_val_1k.txt file provided by Pan et al. (2021) [^3] https://github.com/XingangPan/deep-generative-prior/. For our experiments with the FFHQ dataset, we use the 1000 images (indexed $0,1,\ldots,999$) from the validation set.
E.2 Implementation Details of G2D2 in Inverse Problem Settings
@@
In the continuous optimization phase, we optimize the parameters $\bm{\alpha}$ of the categorical distribution using the RAdam optimizer (Liu et al., 2020). The optimization objective is a weighted sum of the KL divergence term and the likelihood term, defined as:
@@
The marginalization over $\mathbf{z}_{0}$ in line 10 of Algorithm 1, specifically the term $\sum_{\mathbf{z}_{0}}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$, can be computed in closed form. This computation is feasible because both distributions involved in the marginalization are dimensionally independent categorical distributions, as discussed by Austin et al. (2021) and Gu et al. (2022).
@@
Some parameters are dynamically adjusted during inference. Both the learning rate for RAdam ($l_{\text{RAdam}}$) and the KL divergence coefficient ($\eta_{\text{KL}}$) are scheduled using weight vectors that decay logarithmically over the inference steps. These weights are computed based on initial scaling factors.
@@
$l_{\text{RAdam}}(t)=l_{\text{RAdam, base}}\cdot w_{\text{lr}}(t),\quad\eta_{\text{KL}}(t)=\eta_{\text{KL, base}}\cdot w_{\text{KL}}(t).$
@@
$l_{\text{RAdam, base}}$ 15.0
$\lambda_{\text{lr, schedule}}$ 1.0
ImageNet Super-resolution $\eta_{\text{KL, base}}$ 0.0003
@@
$l_{\text{RAdam, base}}$ 10.0
$\lambda_{\text{lr, schedule}}$ 1.0
@@
$l_{\text{RAdam, base}}$ 15.0
$\lambda_{\text{lr, schedule}}$ 1.0
@@
$l_{\text{RAdam, base}}$ 10.0
@@
E.3 G2D2 with Markov Noise Process
@@
We compare the computational complexity of G2D2 with the star-shaped noise process (Algorithm 1) and its hypothetical Markovian counterpart (Algorithm 2). The main algorithmic difference lies in line 10 of both algorithms: the sampling of $\mathbf{z}_{t-1}$ given the optimized $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$.
The key differences between both algorithms are as follows. In G2D2 (Algorithm 1), we have $p_{\alpha}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})=\sum_{\mathbf{z}_{0}}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$, where $q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})$ represents the direct transition from $\mathbf{z}_{0}$ to $\mathbf{z}_{t-1}$ based on the cumulative transition matrix $\bar{Q}_{t-1}$ (derived from $Q_{1},\dots,Q_{t-1}$ as described in Appendix E.2). In the Markovian Version (Algorithm 2), we have $p_{\alpha}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})=\sum_{\mathbf{z}_{0}}q_{\text{Markov}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t})\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$, where $q_{\text{Markov}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t})$ is the posterior from the original Markov process in (35).
Crucially, the computational cost of sampling $\mathbf{z}_{t-1}$ in both cases is dominated by the marginalization over $\mathbf{z}_{0}$ after $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ has been optimized. The calculation of $q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})$ and $q_{\text{Markov}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t})$ involves operations on categorical distributions and transition matrices, which are generally efficient and do not introduce significant computational differences between the two approaches.
@@
5:    Initialize: $\bm{\alpha}_{t}\propto\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$
@@
7:    Initialize: $\bm{\alpha}_{t}\propto\exp(\gamma\log\bm{\alpha}_{t+1}+(1-\gamma)\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t}))$
@@
E.4 Settings for comparison methods
In this subsection, we detail the experimental settings for the comparison methods.
@@ E.4 Settings for comparison methods
(Chung et al., 2023b) We use the same parameter settings as described in the original paper. The guidance scale is set to 1.0 for FFHQ & super-resolution, 1.0 for FFHQ & Gaussian deblurring, 1.0 for ImageNet & super-resolution, and 0.4 for ImageNet & Gaussian deblurring. The number of time steps is set to 1000. For pre-trained models, we use the unconditional model provided by Dhariwal & Nichol (2021)  [^5] https://github.com/openai/guided-diffusion for ImageNet. For FFHQ, we use the model provided by Choi et al. (2021)  [^6] https://github.com/jychoi118/ilvr_adm.
@@
(Rout et al., 2023) We use the official implementation  [^8] https://github.com/LituRout/PSLD. For the pre-trained model, we employ stable-diffusion v-1.5 (Rombach et al., 2022)  [^9] https://github.com/CompVis/stable-diffusion. As this model handles $\times512$ pixel images, we first upscale the ground truth image to $\times512$. We then apply the forward operator to the upscaled image and use the result as observed data for our method. Finally, we downsample the output to $\times256$. For hyperparameters, we use $\eta=1.0$ and $\gamma=0.1$.
@@
E.5 GPU memory usage and computational speed
We analyze the GPU memory consumption and computational speed of our proposed method, G2D2, in comparison with other methods. Table 5 presents an overview of these metrics for various methods. The measurements are conducted using a single NVIDIA A6000 GPU for the Gaussian deblurring task on ImageNet. G2D2 has the lowest memory usage among all methods and the fastest computational speed among gradient-based methods.
@@ E.5 GPU memory usage and computational speed
Method GPU Memory Usage (GiB) Wall-Clock time (s)
@@
E.6 Impact of the Forget Coefficient
@@
E.7 Impact of Text Conditioning on the Prior Model
To examine the necessity of text conditioning, we investigate the effect of the presence or absence of prompts given to VQ-Diffusion on performance. Table 6 shows the performance for each setting. “Not Used” for text conditioning indicates that classifier-free guidance in the prior model is set to 1.0 (equivalent to unconditional sampling). The prompts we provide to VQ-Diffusion in our method are “a photo of [Class Name]” for ImageNet experiments and “a high-quality headshot of a person” for FFHQ experiments. It should be noted that these prompts are extremely general and do not describe specific details of the images.
From these results, we can confirm that prompt conditioning contributes to a certain level of performance improvement on ImageNet, while it does not significantly affect performance on FFHQ. This suggests that the pre-trained VQ-Diffusion model may not have been extensively trained on human face images, or that our chosen prompts for FFHQ may not be optimal.
Additionally, Figure 7 shows a qualitative comparison for the Gaussian deblurring task on the ImageNet dataset. When prompts are not used, smoother results are obtained throughout the intermediate steps when compared, demonstrating that the presence of text conditioning leads to final results that capture more fine-grained details.
Figure 7: Images sampled from the prior model $\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ using intermediate $\mathbf{z}_{t}$ during the process of G2D2 in image inverse problem solving. The top two rows show intermediate samples using the prompt “a photo of guenon, guenon monkey,” while the bottom two rows show results without using any prompt. Each column represents a different time step in the G2D2 algorithm, ranging from $t=95$ (earliest stage) to $t=5$ (final stage). The rightmost column displays the measurement (degraded input image) and ground truth (original image).
Table 6: Performance comparison with and without text conditioning
Dataset Text conditioning SR ($\times 4$) Gaussian Deblurring
[…]
ImageNet Not Used 0.357 23.55 0.385 23.24
Used 0.351 23.37 0.370 23.14
FFHQ Not Used 0.258 27.48 0.274 26.99
Used 0.259 27.44 0.273 27.00
E.8 Impact of prior selection for G2D2
G2D2 can also use other prior models when solving inverse problems. In this paper, we primarily use VQ-Diffusion trained on the ITHQ dataset (hereafter VQDiff-ITHQ) as a prior model, but to investigate how different prior models affect inverse problem solving performance, we conduct experiments using another model, CLIP-VQDiffusion (Han & Kim, 2024)  [^12] https://github.com/INFINIQ-AI1/CLIPVQDiffusion, trained on the FFHQ dataset (hereafter CVQDiff-FFHQ). As in the main text, we evaluated on Gaussian deblurring tasks and super-resolution tasks (with measurement noise $\sigma=0.05$). For evaluation, we used 100 images from the FFHQ validation subset and measured PSNR, LPIPS.
Table 7: Performance comparison between different prior models on FFHQ dataset
Dataset Prior Model SR ($\times 4$) Gaussian Deblurring
[…]
FFHQ VQDiff-ITHQ (Tang et al., 2022) 27.44 0.259 27.00 0.273
CVQDiff-FFHQ (Han & Kim, 2024) 25.32 0.283 24.41 0.289
Figure 8: Qualitative comparison of super-resolution results ($\times$4, $\sigma=0.05$) on FFHQ test images. From left to right: low-resolution measurement, ground truth, G2D2 with VQDiff-ITHQ prior, and G2D2 with CVQDiff-FFHQ prior. While both prior models produce visually pleasing reconstructions, VQDiff-ITHQ prior tends to generate results closer to the ground truth, whereas CVQDiff-FFHQ prior produces fewer artifacts but with slightly less fidelity to the ground truth images.
Table 7 shows the quantitative results, and Figure 8 presents examples of reconstructed samples. Interestingly, although CVQDiff-FFHQ prior is trained on the FFHQ dataset, G2D2 with VQDiff-ITHQ prior performs slightly better. When observing qualitative results, G2D2 with CVQDiff-FFHQ prior produces fewer artifacts but seems to deviate slightly from the ground truth images. Since CVQDiff-FFHQ prior is trained on FFHQ, it likely has an advantage in prior modeling (corresponding to the first term of the optimization target in Algorithm 1, L10). However, the optimization of the likelihood in the measurement equation (the second term) involves decoder calculations, and since both models use different decoders, this difference may be influencing the results.
E.9 Failure modes of G2D2
We conduct an analysis of failure modes. Figure 9 shows the results of G2D2 and the images during inference for the Gaussian deblurring task on FFHQ. When the ground truth image is a relatively young (child’s) face, the generated face images appear to be drawn towards a distribution of more adult faces. This is likely due to the use of the prompt “a high-quality headshot of a person”. As a result, there is a consistent bias towards adult face images throughout the generation process, leading to artifacts in the final image. In the absence of a prompt, the intermediate generated images are not influenced by any specific textual guidance. As a result, the final image tends to have fewer artifacts.
While the star-shaped noise process can correct early errors, if errors persist until the later stages, it becomes more difficult to correct them from that point onwards. In other words, when there is a mismatch between the distribution conditioned by the prompt and the target image, it becomes challenging for G2D2 to handle it effectively.
To improve these issues, techniques such as simultaneous optimization of prompts may be necessary. Prompt-tuning techniques, as proposed in reference (Chung et al., 2024), could be effective in addressing these challenges.
Figure 9: Failure modes of G2D2: Gaussian Deblurring results on the FFHQ dataset. Due to the mismatch between the prompt and the target image, errors remain uncorrected throughout the process, resulting in artifacts in the estimated image.
E.10 Additional quantitative results of G2D2 and comparison methods.
Tables 8-11 shows the evaluation results with SSIM, FID, and KID added as evaluation metrics.
Tables 811 show the evaluation results with Structural Similarity Index Measure (SSIM) (Wang et al., 2004), Fréchet Inception Distance (FID) (Heusel et al., 2017), and Kernel Inception Distance (KID) (Bińkowski et al., 2018) added as evaluation metrics. The evaluation of FID and KID in this study is conducted under the following conditions: We use the clean-fid [^13] https://github.com/GaParmar/clean-fid library for our evaluation. Following the default settings of the clean-fid library, we use 2048-dimensional features extracted from the pool_3 layer of the Inception V3 model. We follow the internal processing of the clean-fid library for image preprocessing, which typically includes resizing input images to $\times299$ pixels and appropriate normalization. For reference statistics, we use pre-computed statistics of the trainval split of the FFHQ dataset (resolution $\times256$, based on 50,000 images) provided by the clean-fid library (specified as dataset_name=‘‘ffhq’’, dataset_res=256, dataset_split=‘‘trainval’’). For the ImageNet dataset ($\times256$ resolution), it should be noted that the reference set for FID calculation on ImageNet consists of 1,000 images, which is smaller than the recommended sample size for typical FID evaluation (e.g., 10,000-50,000 images). Therefore, FID scores under these conditions should be interpreted with consideration of the limitations regarding the reference set size.
Table 8: Quantitative evaluation on ImageNet $\times256$. Performance comparison of different methods on Super-Resolution ($\times$4) task. Values show mean over 1000 images.
Prior Type Method PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) FID($\downarrow$) KID($\downarrow$)($\times 10^{-2}$)
Pixel-domain DPS 22.67 0.362 0.618 45.97 0.41
DDRM 24.27 0.351 0.669 52.26 1.16
LDM PSLD 24.02 0.331 0.675 47.88 0.83
ReSample 23.31 0.373 0.616 68.77 1.96
Discrete G2D2 (proposed) 23.82 0.340 0.638 49.97 0.86
G2D2 w/ Markov noise process 22.01 0.442 0.557 78.43 2.40
Table 9: Quantitative evaluation on ImageNet $\times256$. Performance comparison of different methods on Gaussian Deblurring task. Values show mean over 1000 images.
Prior Type Method PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) FID($\downarrow$) KID($\downarrow$)($\times 10^{-2}$)
Pixel-domain DPS 19.49 0.432 0.473 62.04 0.86
DDRM 27.61 0.250 0.786 34.05 0.40
LDM PSLD 23.95 0.359 0.659 50.96 0.96
ReSample 23.07 0.425 0.586 82.57 2.44
Discrete G2D2 (proposed) 23.37 0.367 0.604 61.60 1.39
G2D2 w/ Markov noise process 22.36 0.424 0.551 79.29 2.43
Table 10: Quantitative evaluation on FFHQ $\times256$. Performance comparison of different methods on Super-Resolution ($\times$4) task. Values show the mean over 1000 images.
Prior Type Method PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) FID($\downarrow$) KID($\downarrow$)($\times 10^{-2}$)
Pixel-domain DPS 26.07 0.238 0.756 26.85 0.67
DDRM 28.09 0.252 0.804 46.19 3.00
LDM PSLD 27.12 0.282 0.757 37.37 2.07
ReSample 23.07 0.508 0.445 85.86 7.05
Discrete G2D2 (proposed) 27.29 0.265 0.763 45.21 3.08
G2D2 w/ Markov noise process 25.15 0.369 0.699 62.96 4.47
Table 11: Quantitative evaluation on FFHQ $\times256$. Performance comparison of different methods on Gaussian Deblurring task. Values show the mean over 1000 images.
Prior Type Method PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) FID($\downarrow$) KID($\downarrow$)($\times 10^{-2}$)
Pixel-domain DPS 25.47 0.234 0.731 24.10 0.40
DDRM 30.89 0.209 0.863 42.51 2.51
LDM PSLD 26.96 0.307 0.750 38.84 1.99
ReSample 25.91 0.336 0.642 45.26 2.85
Discrete G2D2 (proposed) 26.91 0.280 0.743 47.70 3.46
G2D2 w/ Markov noise process 25.96 0.340 0.708 57.73 4.29
E.11 Additional qualitative results of G2D2 and comparison methods.